If w represents the width of a rectangle 12 inches long with an area of 500 square inches, what is the best algebraic equation for the area of the rectangle?

Answer by Alan3354(21583)

If w represents the width of a rectangle 12 inches long with an area of 500 square inches, what is the best algebraic equation for the area of the rectangle?
--------------
The equation for the area of any rectangle is length x width, or
A = L*W
For this one,
500 = 12L
L = 500/12 = 125/3 inches
There's no reason to deal with the equation once the measurements are known.